If m = 2, find the value of:
5m/2 – 4 …...................

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

From the question it is given that m = 2
Then, substitute the value of p in the question
= ((5 × 2)/2)–4
= (10/2)–4
= 5–4
= 1

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Given : m = 2
Substituting value of m in the equation,
5m/2 - 4 
= (5×2)/2 - 4
= 10/2 - 4
= 5-4
= 1
 
Hope it helps!
Thank you ✌️